OpenStudy (anonymous):

Students who work an average of at least 16hr per week receive extra pay and part time employee benefits. for the first three weeks of september, dina worked 20hr,12hr, and 14hr. how many hours must she work during the fourth week in order to average at least 16 hr per week for the month?

OpenStudy (shane_b):

You just need to find the missing value in the following:\[\frac{20+12+14+x}{4}=16\]

OpenStudy (shane_b):

20+12+14 = 36 so you can start by simplifying and splitting up the fraction:\[\frac{36}{4}+\frac{x}{4}=16\]\[\frac{36}{4}=9\] Now subtract 9 from both sides to get:\[\frac{x}{4}=7\]Multiply both sides by 4 to get:\[(\cancel{4})\frac{x}{\cancel{4}}=7(4)\]Therefore, x = 28